# Polaris: Sampling from the Multigraph Configuration Model with Prescribed Color Assortativity

### v1

@@ 2. Related Work
When searching for patterns in network data it is essential to be able to reason about their significance. In statistics, there is a long tradition of assessing significance by comparing an observed pattern with its occurrence in a randomized null model (Fisher, 1935). Extending this idea to networks leads to random-graph null models, where one compares properties observed in real-world networks with properties observed in networks sampled from a certain random-graph distribution. While there is a vast literature on random-graph models, such as the Erdős-Rényi random graph (Erdős and Rényi, 1959), the preferential attachment model (Barabási and Albert, 1999; Bollobás et al., 2001), and many other, which are simple to generate via an iterative sampling process, practitioners are often interested in sampling networks from a space of networks satisfying certain constraints. The most commonly-used constrained random-graph null model is the configuration model (Bender and Canfield, 1978; Bollobás, 1980; Fosdick et al., 2017), where the sample space consists of all networks having a specified degree sequence. Configuration models have a long history in research and found many applications in sociology (Moreno and Jennings, 1938), ecology (Connor and Simberloff, 1979), systems biology (Milo et al., 2002), and other disciplines.
@@ 3.2. Markov Chain Monte Carlo Methods
All algorithms in Polaris- $\ast$ follow the Markov chain Monte Carlo (MCMC) method, using the Metropolis-Hastings (MH) approach (Mitzenmacher and Upfal, 2005, Ch. 7 and 10). We now introduce these concepts, tailoring them to our purposes.
Let $\mathcal{G}=(\mathcal{S},\mathcal{E},\mathsf{w})$ be a directed, weighted, strongly connected, and aperiodic [^1] A graph is aperiodic iff the greatest common denominator of the lengths of its cycles is 1. graph, which may have self-loops. The vertices in $\mathcal{S}$ are known as states, and $\mathcal{G}$ is known as a state graph. Using the same notation we defined previously, for any state $s\in\mathcal{S}$, $\Gamma_{\mathcal{G}}(s)$ denotes the set of (out-)neighbors of $s$, i.e., the set of states $u$ s.t. $(s,u)\in\mathcal{E}$. Iff $u\in\Gamma_{\mathcal{G}}(s)$, then $\mathsf{w}(s,u)>0$, and it holds $\sum_{u\in\Gamma_{\mathcal{G}}(s)}\mathsf{w}(s,u)=1$. Thus, for any $u\in\mathcal{S}$, we can define the transition probability $\tau_{s,u}$ from $s$ to $u$ as $\mathsf{w}(s,u)$ if $u\in\Gamma_{\mathcal{G}}(s)$, and 0 otherwise.
@@ 4. A Null Model for Vertex-Colored Graphs
Given an observed $\mathring{G}\doteq(V,E,\mathcal{L},\lambda)$, with $V=\{v_{1},\dotsc,v_{\left\lvert V\right\rvert}\}$, we consider the null model $\Pi=(\mathcal{Z},\pi)$ where $\mathcal{P}_{\mathring{G}}$ consists of the following properties:
  • P1:
    degree sequence $\left[d_{\mathring{G}}(v_{1}),\dotsc,d_{\mathring{G}}\left(v_{\left\lvert V\right\rvert}\right)\right]$;
  • P2:
    JCM $J_{\mathring{G}}$.
@@ 4.1. Strong Connectivity and Aperiodicity of $\mathcal{Z}$ via JDESs
The proof of the above theorem explicitly builds a sequence of JDESs from any state $G\in\mathcal{Z}$ to any other $H\in\mathcal{Z}$, by first going from $G$ to a $\tilde{G}\in\mathcal{Z}$ such that every vertex has in $\tilde{G}$ exactly the same number of neighbors of each color as it has in $H$, and then going from $\tilde{G}$ to $H$.
@@
Given a multigraph $G$, if either of the following conditions holds:
@@ 4.3. An algorithm tailored to the task
    if the sampled pair of distinct edges is such that one of the DESs involving is a no-op or not a JDES and the other is a moving JDES (Cases 2B, 2C, 3B, 3C) Polaris-C always chooses the moving one;
@@
    if the sampled pair of distinct edges is such that the JDESs involving them are equivalent (Cases 1, 2A, 2D, 3A), it avoids sampling one of them at random, and deterministically chooses one.
@@ 5. Experimental evaluation
Our experimental evaluation has three objectives. First, we demonstrate the qualitative differences between multigraphs sampled using the traditional configuration model and those obtained from Polaris. Second, we analyze the extent to which the baseline algorithm Polaris-B differs from the color-aware algorithm Polaris-C in their respective movements within the state space. Lastly, we show the scalability of both Polaris-B and Polaris-C, particularly in relation to the number of vertex colors and the number of edges.
The code and the datasets used are available on GitHub. [^3] https://github.com/lady-bluecopper/Polaris
Cite 3264 4611 6 2.83 2.00 0.17 0.18
Brexit 22745 48830 2 4.29 1.00 0.50 0.50
Twitter 22405 77920 3 6.96 1.00 0.33 0.32
Phy-Cit 30501 347268 11 22.77 14.00 0.09 0.11
Abortion 279505 671144 2 4.80 1.00 0.50 0.50
US-Elect 23832 845152 3 70.93 3.00 0.33 0.25
Trivago 172738 1327092 160 15.37 6.00 0.01 0.00
Obamacare 334617 1511670 2 9.04 1.00 0.50 0.50
Walmart 88860 2267396 11 51.03 18.00 0.09 0.05
Comb 677753 6666018 2 19.67 1.00 0.50 0.50
Guns 632659 7478993 2 23.64 1.00 0.50 0.50
Datasets. We consider $11$ real-world labeled networks, whose characteristics are summarized in Table 1. Brexit, US-Elect, Abortion (Garimella et al., 2017b), Twitter (Conover et al., 2011), Obamacare (Garimella et al., 2017c), Comb, and Guns (Garimella et al., 2018a) are retweet networks generated from tweets collected on various controversial topics. An edge exists between two users if one retweeted the other. Node colors indicate the side taken in the discussion, with a third label indicating neutrality. Cite and Phy-Cit  (Preti et al., 2023) are citation networks: nodes represent publications, with node colors indicating Computer Science areas and the year of publication, respectively. Trivago (Chodrow et al., 2021) is network where nodes are accommodations, and edges connect accommodations visited by a user in the same browsing session. Node colors indicate the country where the accommodation is located. Walmart (Amburg et al., 2020) is a co-purchase network where nodes are Walmart products, and edges connect products that were bought together. Node colors indicate the departments in which the products appear on walmart.com.
Experimental Setup. All experiments are run on an Intel Xeon Silver 4210R CPU@2.40GHz running FreeBSD with $383$ GiB of RAM. We evaluate three sampling algorithms: the baseline color-agnostic algorithm Polaris-B, the color-aware algorithm Polaris-C, and the traditional configuration model (CM), which samples from the state space of multigraphs with a prescribed degree sequence.
@@ 5.1. Comparison with the Configuration Model
Figure 2 shows that the color assortativity values of the multigraphs sampled by CM significantly diverge from those of the corresponding observed multigraphs, with relative errors close to $1$. This discrepancy is more pronounced in datasets with a larger number of colors. This result proves that CM does not adequately capture the color assortativity present in the observed data.
@@
Figure 4 presents the running time required by Polaris-B and Polaris-C to perform $m\log(m)$ iterations on different versions of the Walmart dataset (left), and the distribution of the relative errors of color assortativity of the samples generated by CM (right). Starting from the $11$ available colors (corresponding to product categories), we cluster these colors to create new realistic sets of $2$, $4$, and $8$ colors.
@@ 6. Conclusion
We introduced Polaris, an ensemble of colored multigraphs with prescribed Joint Color Matrix. The JCM captures important properties that are interesting in the study of polarized networks, such as color assortativity. We described two efficient algorithms to sample from the space of such multigraphs, according to any user-specified probability distribution over this space. Our algorithms work by running a Markov chain on the multigraph space, following the Metropolis-Hastings approach for determining whether to accept the move to a proposed neighbor of the current state. We conduct an extensive experimental evaluation, showing the shortcomings of existing methods in capturing the color assortativity, and assessing the performance of our algorithms on different datasets in terms of scalability, runtime, and acceptance probability.
@@ A.1. Proofs
If the edges fall in Cases 2A or 3A, then applying either of the equivalent JDESs involving the edges leads to a state $H\neq G$, and in this state, the new edges fall in case 2B or either 3B or 3C respectively, meaning that the state graph has a self-loop on $H$, thus it is aperiodic.
If the edges fall in case 4C, [^4] The proof for this case is the same as the proof for the aperiodicity of the state graph by DESs (Fosdick et al., 2017, Lemma 3). the state graph has a cycle of length two, because every JDES is reversible. It also has a cycle of length three, by applying the following sequence of JDESs: $(u,w),(v,z)\rightarrow(u,v),(w,z)$, $(u,v),(w,z)\rightarrow(u,z),(v,w)$, $(u,z),(v,w)\rightarrow(u,w),(v,z)$. The greatest common divisor of the lengths of these two cycles is one, thus the state graph is aperiodic.
@@
Let $G,H\in\mathcal{Z}$, $H\neq G$. For any $\ell\in\mathcal{L}$, it holds $\left\lvert E_{G,\ell}\right\rvert=\left\lvert E_{H,\ell}\right\rvert$, i.e., the size of these sets is constant across all multigraphs in $\mathcal{Z}$.
@@
A.2. State Space is Disconnected for Simple Graphs with More Than 2 Colors
We first redefine the JDES for the case of simple graphs.
Definition A.5 (JDES).
Given a graph $G\doteq(V,E,\mathcal{L},\lambda)$, let $(u,w),$ $(v,z)\in E$ such that $u\neq v$, $w\neq z$, $(u,z)\notin E$, $(v,w)\notin E$, and $\lambda(u)=\lambda(v)$. Consider the graph $G^{\prime}=(V,E\setminus\{(u,w),(v,z)\}\cup\{(u,z),(w,v)\},\mathcal{L},\lambda)$. Clearly, $G^{\prime}$ has the same degree sequence and JCM as $G$.
We now show a counterexample proving that JDESs cannot strongly connect the state space $\mathcal{Z}$ of graphs with the same degree distribution and JCM when $\left\lvert\mathcal{L}\right\rvert>2$, i.e., when there are more than two colors.
Let $k>2$ and =$\mathcal{L}=\{c_{1},\dotsc,c_{k}\}$. Let $V=\{1,\dotsc,2k\}$ be a set of vertices with $\lambda(i)=\lambda(k+i)=c_{i}$, $1\leq i\leq k$. We now build two graphs $G^{\prime}=(V,E^{\prime},\mathcal{L},\lambda)$ and $G^{\prime\prime}=(V,E^{\prime\prime},\mathcal{L},\lambda)$ with the same degree sequence and the same JCM, but for which there is no sequence of JDESs transforming $G^{\prime}$ to $G^{\prime\prime}$.
In both graphs, vertices $1,\dotsc,k$ form a clique. In addition to these edges, $G^{\prime}$ contains the $k$ edges $\{(1,k+2),(2,k+3),\dotsc,(k-1,2k),(k,k+1)\}$, while $G^{\prime\prime}$ contains the $k$ edges $\{(1,2k),(2,k+1),(3,k+2),\dotsc,(k,2k-1)\}$.
The two graphs have the same degree sequence and the same JCM, thus they both belong to $\mathcal{Z}$, taking either of them to be $\mathring{G}$. Figure 7 shows $G^{\prime}$ (top) and $G^{\prime\prime}$ (bottom).
Figure 7. The graphs $G^{\prime}$ (upper) and $G^{\prime\prime}$ (lower) used in the counterexample.
[…]
Drawings of the graphs $G^{\prime}$ and $G^{\prime\prime}$ used in the counterexample.
The fact that $\mathcal{Z}$ is not strongly-connected by JDESs is a corollary of the following lemma.
Lemma A.6.
There is no JDES from $G^{\prime}$ to any other graph in $\mathcal{Z}$.
Proof.
Assume by contradiction that there exists a JDES $(u,w),$ $(v,z)$rightarrow $(u,z),(v,w)$ from $G^{\prime}$. According to Definition A.5, it must be the case that $u=i$ and $v=k+i$ for some $1\leq i\leq k$, as it must be $\lambda(u)=\lambda(v)$. W.l.o.g., assume that $u=1$ and $v=k+1$. Let $z_{i}$ be the element in $[1,k]$ s.t. $z_{i}\equiv(k+i-1)\mod k$. Since $k+1$ has degree $1$, we must have $z=z_{i}$. Regardless of the choice of $w$, the edge swap $(u,w),(v,z)\rightarrow(u,z),(v,w)$ would result in the creation of the multi-edge $(1,k)$, which violates the conditions for a JDES. Thus, it is not possible to construct any JDES from $G^{\prime}$. ∎
In this particular case, $G^{\prime}$ and $G^{\prime\prime}$ can be connected via $k$-switch operations (Tabourier et al., 2011), which are edge swap operations involving the simultaneous swap of $k$ edges. Specifically, the $k$-switch operation involving the edge list $[(1,k+2),(2,k+3),\dotsc,(k-1,2k),(k,k+1)]$, where the $i$-th edge is swapped with the $(i-2)$-th edge for $i\in[3,k]$, the first edge with the $(k-1)$-th edge, and the second edge with the $k$-th edge, allows the transformation of $G^{\prime}$ into $G^{\prime\prime}$. However, this fact alone does not imply that the state space is strongly connected by $k$-switches. Investigating whether the state space is indeed strongly connected under $k$-switches, and addressing the associated computational challenges are open questions for future research.

### v2

@@ Abstract.
Hypothesis Testing, Null Model, Polarization
ccs: Information systems Web miningccs: Theory of computation Graph algorithms analysisccs: Theory of computation Random walks and Markov chainsccs: Theory of computation Generating random combinatorial structuresccs: Mathematics of computing Random graphsccs: Theory of computation Social networks
@@ 2. Related Work
When searching for patterns in network data it is essential to be able to reason about their significance. In statistics, there is a long tradition of assessing significance by comparing an observed pattern with its occurrence in a randomized null model (Fisher, 1935). Extending this idea to networks leads to random-graph null models, where one compares properties observed in real-world networks with properties observed in networks sampled from a certain random-graph distribution. While there is a vast literature on random-graph models, such as the Erdős-Rényi random graph (Erdős and Rényi, 1959) and the preferential attachment model (Barabási and Albert, 1999; Bollobás et al., 2001), which are simple to generate via an iterative sampling process, practitioners often seek to sample networks from a space of networks satisfying certain constraints. The most commonly-used constrained random-graph null model is the configuration model (Bender and Canfield, 1978; Bollobás, 1980; Fosdick et al., 2017), where the sample space consists of all networks having a specified degree sequence. Configuration models have a long research history with applications in sociology (Moreno and Jennings, 1938), ecology (Connor and Simberloff, 1979), systems biology (Milo et al., 2002), and other disciplines.
@@ 3.2. Markov Chain Monte Carlo Methods
All algorithms in Polaris- $\ast$ follow the Markov chain Monte Carlo (MCMC) method, using the Metropolis-Hastings (MH) approach (Mitzenmacher and Upfal, 2005, Ch. 7 and 10). Let us now introduce these concepts.
Let $\mathcal{G}=(\mathcal{S},\mathcal{E},\mathsf{w})$ be a directed, weighted, strongly connected, and aperiodic [^1] A graph is aperiodic iff the greatest common divisor of the lengths of its cycles is 1. graph, which may have self-loops. The vertices in $\mathcal{S}$ are known as states, and $\mathcal{G}$ is known as a state graph. Using the same notation we defined previously, for any state $s\in\mathcal{S}$, $\Gamma_{\mathcal{G}}(s)$ denotes the set of (out-)neighbors of $s$, i.e., the set of states $u$ s.t. $(s,u)\in\mathcal{E}$. Iff $u\in\Gamma_{\mathcal{G}}(s)$, then $\mathsf{w}(s,u)>0$, and it holds $\sum_{u\in\Gamma_{\mathcal{G}}(s)}\mathsf{w}(s,u)=1$. Thus, for any $u\in\mathcal{S}$, we can define the transition probability $\tau_{s,u}$ from $s$ to $u$ as $\mathsf{w}(s,u)$ if $u\in\Gamma_{\mathcal{G}}(s)$, and 0 otherwise.
@@ 4. A Null Model for Vertex-Colored Graphs
Given an observed $\mathring{G}\doteq(V,E,\mathcal{L},\lambda)$, with $V=\{v_{1},\dotsc,v_{\left\lvert V\right\rvert}\}$, we consider the null model $\Pi=(\mathcal{Z},\pi)$ where $\mathcal{P}_{\mathring{G}}$ consists of the degree sequence $\left[d_{\mathring{G}}(v_{1}),\dotsc,d_{\mathring{G}}\left(v_{\left\lvert V\right\rvert}\right)\right]$ and the JCM $J_{\mathring{G}}$.
@@ 4.1. Strong Connectivity and Aperiodicity of $\mathcal{Z}$ via JDESs
The proof of this theorem explicitly builds a sequence of JDESs from any state $G\in\mathcal{Z}$ to any other $H\in\mathcal{Z}$, by first going from $G$ to a $\tilde{G}\in\mathcal{Z}$ such that every vertex has in $\tilde{G}$ exactly the same number of neighbors of each color as it has in $H$, then going from $\tilde{G}$ to $H$.
@@
Given a multigraph $G$, if either of the following conditions holds, then the state graph $\mathcal{G}$ is aperiodic:
@@ 4.3. An algorithm tailored to the task
    if the sampled pair of distinct edges is such that one of the DESs is a no-op or not a JDES and the other is a moving JDES (Cases 2B, 2C, 3B, 3C) Polaris-C always chooses the moving one;
@@
    if the sampled pair of distinct edges is such that the JDESs involving them are equivalent (Cases 1, 2A, 2D, 3A), it deterministically chooses one thus avoiding random choices.
@@ 5. Experimental evaluation
Our experimental evaluation has three objectives. First, we demonstrate the qualitative differences between multigraphs sampled using the traditional configuration model and those obtained from Polaris. We focus on the configuration model as it is the standard reference model in network analysis (Newman, 2018) and aligns with the focus of this paper on microcanonical ensembles. Second, we analyze the extent to which the baseline algorithm Polaris-B differs from the color-aware algorithm Polaris-C in their respective movements within the state space. Lastly, we show the scalability of both Polaris-B and Polaris-C, particularly in relation to the number of vertex colors and the number of edges.
Datasets. We consider $11$ real-world labeled networks, whose characteristics are summarized in Table 1 in Section A.2.
Experimental Setup. All experiments are run on an Intel Xeon Silver 4210R CPU@2.40GHz running FreeBSD with $383$ GiB of RAM. We evaluate three sampling algorithms: the baseline color-agnostic algorithm Polaris-B, the color-aware algorithm Polaris-C, and the traditional configuration model (CM), which samples from the state space of multigraphs with a prescribed degree sequence. The code and the datasets used are available on GitHub. [^3] https://github.com/lady-bluecopper/Polaris
@@ 5.1. Comparison with the Configuration Model
Figure 2 shows that the color assortativity values of the multigraphs sampled by CM significantly diverge from those of the corresponding observed multigraphs, with relative errors close to $1$. This discrepancy arises because CM disrupts the original correlations in the observed datasets, generating random graphs with low color assortativity. The effect is more pronounced in datasets with a larger number of colors or higher color assortativity, where the gap between the observed assortativity and that of the randomized graphs is larger. Consequently, we observe larger relative errors in datasets such as Trivago ($|\mathcal{L}|=160$), Obamacare, and Abortion (assortativity 0.95), and smaller errors in datasets such as Comb and Guns ($|\mathcal{L}|=2$ with assortativity values of 0.31 and 0.35, respectively). This result proves that CM does not adequately capture the color assortativity present in the observed data.
@@
Figure 4 presents the running time required by Polaris-B and Polaris-C to perform $m\log(m)$ iterations on different versions of the Walmart dataset (left), and the distribution of the relative errors of color assortativity of the samples generated by CM (right). Starting from the $11$ available colors (product categories), we cluster these colors to create new realistic sets of $2$, $4$, and $8$ colors.
@@ 6. Conclusion
We introduced Polaris, an ensemble of colored multigraphs with prescribed Joint Color Matrix. The JCM captures key properties relevant to the study of polarized networks, such as color assortativity. We described two efficient algorithms to sample from the space of such multigraphs according to any user-specified probability distribution over this space. Our algorithms work by running a Markov chain on the multigraph space, following the Metropolis-Hastings approach for determining whether to accept the move to a proposed neighbor of the current state. We conducted an extensive experimental evaluation, showing the shortcomings of existing methods in capturing the color assortativity, and assessing the performance of our algorithms across different datasets in terms of scalability, runtime, and acceptance probability.
This work serves as an important first step toward analyzing polarization in real networks. While a comprehensive study of polarization lies beyond this paper’s scope, the tools developed here lay the groundwork for future studies on this subject.
@@ 7. Ethical considerations
Acknowledgements.
MR’s work is supported by the Sponsor National Science Foundation https://www.nsf.gov grants IIS-Grant #2006765, and CAREER-Grant #2238693. AG’s work is supported by the ERC Advanced Grant REBOUND (834862), the EC H2020 RIA project SoBigData++ (871042), and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ A.1. Missing Proofs
If the edges fall in Cases 2A or 3A, then applying either of the equivalent JDESs involving the edges leads to a state $H\neq G$ where the new edges fall in case 2B or either 3B or 3C respectively, meaning that the state graph has a self-loop on $H$, thus it is aperiodic.
If the edges fall in case 4C, [^4] The proof for this case is the same as the proof for the aperiodicity of the state graph by DESs (Fosdick et al., 2017, Lemma 3). the state graph has a cycle of length two, because every JDES is reversible. It also has a cycle of length three, by applying the following sequence of JDESs: $(u,w),(v,z)\rightarrow(u,v),(w,z)$, $(u,v),(w,z)\rightarrow(u,z),(v,w)$, and $(u,z),(v,w)\rightarrow(u,w),(v,z)$. The greatest common divisor of the lengths of these two cycles is one, thus the state graph is aperiodic.
@@
Let $G\neq H\in\mathcal{Z}$. For any $\ell\in\mathcal{L}$, it holds $\left\lvert E_{G,\ell}\right\rvert=\left\lvert E_{H,\ell}\right\rvert$, i.e., the size of these sets is constant across all multigraphs in $\mathcal{Z}$.
@@
A.2. Datasets
We consider $11$ real-world labeled networks, whose characteristics are summarized in Table 1. Brexit, US-Elect, Abortion (Garimella et al., 2017b), Twitter (Conover et al., 2011), Obamacare (Garimella et al., 2017c), Comb, and Guns (Garimella et al., 2018a) are retweet networks generated from tweets collected on various controversial topics. An edge exists between two users if one retweeted the other. Node colors indicate the side taken in the discussion, with a third label indicating neutrality. Cite and Phy-Cit  (Preti et al., 2023) are citation networks: nodes represent publications, with node colors indicating Computer Science areas and the year of publication, respectively. Trivago (Chodrow et al., 2021) is network where nodes are accommodations, and edges connect accommodations visited by a user in the same browsing session. Node colors indicate the country where the accommodation is located. Walmart (Amburg et al., 2020) is a co-purchase network where nodes are Walmart products, and edges connect products that were bought together. Node colors indicate the departments in which the products appear on walmart.com.
Cite 3264 4611 6 2.83 2.00 0.17 0.18
Brexit 22745 48830 2 4.29 1.00 0.50 0.50
Twitter 22405 77920 3 6.96 1.00 0.33 0.32
Phy-Cit 30501 347268 11 22.77 14.00 0.09 0.11
Abortion 279505 671144 2 4.80 1.00 0.50 0.50
US-Elect 23832 845152 3 70.93 3.00 0.33 0.25
Trivago 172738 1327092 160 15.37 6.00 0.01 0.00
Obamacare 334617 1511670 2 9.04 1.00 0.50 0.50
Walmart 88860 2267396 11 51.03 18.00 0.09 0.05
Comb 677753 6666018 2 19.67 1.00 0.50 0.50
Guns 632659 7478993 2 23.64 1.00 0.50 0.50
A.3. Computational Complexity
Polaris-C has an initialization step for the data structures used to ensure we only sample pairs of edges that form a JDES. This initialization phase has a time complexity of $\mathcal{O}(n+m)$ and requires $\mathcal{O}(n+m)$ space, where $n$ is the number of vertices and $m$ is the number of edges. The space is used to store information such as node degrees, node colors, the edge list, and the edge subsets $E_{G,\ell}$ for each color $\ell$. Since each edge can belong to up to two of these subsets, they occupy at most $2m$ space in total. Each step of the algorithm then takes constant time, and we perform $s$ steps. In our experiments, we follow previous works and set $s=m\log(m)$ as  (Viger and Latapy, 2005). The number of colors $|\mathcal{L}|$ does not affect the time complexity, as Polaris-C directly samples pairs of edges that can form a JDES. In contrast, Polaris-B is affected by $|\mathcal{L}|$. As $|\mathcal{L}|$ increases, so does the number of possible combinations of node colors, thus reducing the probability that two sampled edges belong to the same subset $E_{G,\ell}$. Each time this condition is not met, Polaris-B must resample two new edges, which increases its running time.